# Story-Adapter: A Training-free Iterative Framework for Long Story Visualization

### v1

@@ Abstract
Story visualization, the task of generating coherent images based on a narrative, has seen significant advancements with the emergence of text-to-image models, particularly diffusion models. However, maintaining semantic consistency, generating high-quality fine-grained interactions, and ensuring computational feasibility remain challenging, especially in long story visualization (i.e., up to 100 frames). In this work, we propose a training-free and computationally efficient framework, termed Story-Adapter, to enhance the generative capability of long stories. Specifically, we propose an iterative paradigm to refine each generated image, leveraging both the text prompt and all generated images from the previous iteration. Central to our framework is a training-free global reference cross-attention module, which aggregates all generated images from the previous iteration to preserve semantic consistency across the entire story, while minimizing computational costs with global embeddings. This iterative process progressively optimizes image generation by repeatedly incorporating text constraints, resulting in more precise and fine-grained interactions. Extensive experiments validate the superiority of Story-Adapter in improving both semantic consistency and generative capability for fine-grained interactions, particularly in long story scenarios. The project page and associated code can be accessed via this https URL.
Figure 1: A long story of “snowman” visualized by our Story-Adapter from different iterations, compared with those visualized by previous StoryDiffusion (Zhou et al., 2024) and StoryGen (Liu et al., 2024). Notable differences are highlighted in green and red. Zoom in for a better view.
@@ 1 Introduction
Story visualization aims to generate a sequence of coherent images from text prompts, reflecting the narrative’s progression and enabling users, even without an artistic background, to visually present their stories (Li et al., 2019; Maharana & Bansal, 2021; Chen et al., 2022). Recent advancements in text-to-image models, particularly diffusion models, have significantly improved the quality of generated visuals, producing high-quality, creative, and aesthetically pleasing images (Saharia et al., 2022; Rombach et al., 2022; Kang et al., 2023). These models greatly outperform earlier approaches such as generative adversarial networks (Brock, 2018) in terms of image quality.
@@
However, story visualization remains challenging, particularly in maintaining semantic consistency and capturing complex interactions as the story length increases. Two main paradigms have emerged in this domain. The Auto-Regressive paradigm (Fig. 2A), which generates frames sequentially (Pan et al., 2024; Liu et al., 2024), often struggles with semantic consistency due to error accumulation and the inability to reference future frames, leading to inconsistencies in the overall narrative. Although techniques like Consistent Self-Attention (CSA) (Zhou et al., 2024) can help mitigate these inconsistencies, their reliance on intermediate denoising features results in high memory consumption, limiting scalability for longer stories. To address these challenges, Zhou et al. (2024) further propose the Reference-Image paradigm, which employs fixed reference images to guide the visualization process. However, as shown in Fig. 2B, while using only the initial frames as reference images alleviate scalability issues, it fails to provide the global semantic coherence necessary for long-story visualization, ultimately resulting in the propagation of errors from the reference images to subsequent frames. As such, both paradigms experience quality degradation when visualizing long stories. Additionally, they inherit the limitations from Stable Diffusion (SD) (Rombach et al., 2022), particularly in generating fine-grained interactions (as shown in Fig. 1).
To address these limitations, we present Story-Adapter, an iterative framework that adapts pre-trained SD models for long story visualization. Unlike existing methods that generate images auto-regressively or rely on static reference images (Fig. 2 A&B), our approach prioritizes semantic consistency by incorporating all generated images from previous iterations into the current one. This process offers two key advantages. 1) It offers a comprehensive view of the entire narrative, thereby reducing error accumulation and mitigating the propagation of flaws from reference images. 2) By continuously engaging with text prompts, Story-Adapter optimizes generative quality for details based on insights from earlier iterations. As illustrated in Fig.1, our framework enhances both semantic consistency and the quality of fine-grained interactions across iterations, resulting in more coherent and higher-quality visualizations. For example, the image depicting complex character interactions, such as “the snowman greeting the fox” demonstrates substantial improvement over iterations compared to previous methods(Liu et al., 2024; Zhou et al., 2024).
@@
During initialization, only text prompts of the story are utilized to generate reference images. In subsequent iterations, the global embeddings of all images generated in the previous round, along with the text embeddings, collaboratively guide the image generation process. To implement the iterative paradigm efficiently, we propose a plug-and-play Global Reference Cross-Attention (GRCA), where all global image embeddings act as keys and values. This significantly reduces computational costs, as global embeddings operate at a lower dimensionality than the intermediate denoising features used in CSA. Additionally, to strike a balance between visual consistency and text controllability, we introduce a linear weighting strategy in the iterative paradigm to fuse both modalities.
@@
Extensive experiments demonstrate that Story-Adapter consistently outperforms existing methods for visualizing both regular-length and long stories (up to 100 frames). Specifically, in the context of regular-length story visualization using the StorySalon benchmark dataset (Liu et al., 2024), Story-Adapter exceeds the baseline model, StoryGen (Liu et al., 2024), achieving a 9.4% improvement in average Character-Character Similarity (aCCS) (Cheng et al., 2024) and a 21.71 reduction in average Fréchet Inception Distance (aFID) (Cheng et al., 2024). For long story visualization, Story-Adapter also demonstrates solid advancements, achieving gains of 3.4% in aCCS and 8.14 in aFID compared to StoryDiffusion (Zhou et al., 2024), demonstrating the superior generative quality of Story-Adapter, particularly in terms of semantic consistency and fine-grained interactions.
Figure 2: Comparison of paradigms for long story visualization: (A) Auto-Regressive (AR): generates frames sequentially referencing on previous finite frames (e.g. the previous three frames); (B) Reference-Image (RI): employs fixed reference images (e.g. the beginning four frames) as reference images; (C) Iterative Paradigm: leverages all frames from the previous iteration as reference images.
@@ 2 Related Work
2.1 Diffusion Models
Diffusion models (Ho et al., 2020; Song et al., 2020b; Sohl-Dickstein et al., 2015; Nichol & Dhariwal, 2021) have emerged as powerful tools for data distribution modeling through iterative denoising. Recent advancements in sampling techniques (Xiao et al., 2021; Song et al., 2020a; Luo et al., 2023), backbone architectures (Peebles & Xie, 2023; Lu et al., 2023), and latent space denoising (Rombach et al., 2022; Podell et al., 2023) have led to their widespread adoption in various generative tasks, including video (Esser et al., 2023; Yang et al., 2024), 3D (Luo & Hu, 2021; Xu et al., 2024), audio (Ruan et al., 2023; Huang et al., 2023), and human motion generation (Zhang et al., 2022; Karunratanakul et al., 2023). While Text-to-Image diffusion models (Saharia et al., 2022; Zhang et al., 2023; Rombach et al., 2022; Podell et al., 2023) have gained significant attention, challenges persist in generating coherent image sequences for tasks like story visualization due to the inherent randomness and fine-grained interaction generation.
2.2 Story Visualization
Story visualization (Chen et al., 2022; Li, 2022) has evolved from GAN-based approaches like StoryGAN (Li et al., 2019) to more advanced techniques. Recent developments leverage diffusion models (Shen et al., 2024; Tao et al., 2024) and combine them with auto-regressive paradigm, as seen in AR-LDM (Pan et al., 2024) and StoryGen (Liu et al., 2024). These methods have improved coherence in image sequences and extended to open-ended story visualization. However, challenges remain in maintaining semantic consistency for the whole story and avoiding error accumulation, especially for longer narratives (Wang et al., 2023; Zhou et al., 2024; Liu et al., 2024).
2.3 Subject-Consistent Image Generation
The consistency of the generated subjects is critical for tasks such as story visualization and video generation. Recent advancements in subject-consistent image generation have focused on reducing computational resources while maintaining consistency. Early approaches like Gal et al. (2022); Ruiz et al. (2023) require extensive fine-tuning, prompting more efficient methods (Ryu, 2023; Han et al., 2023; Kumari et al., 2023; Yuan et al., 2023). Notable progress includes IP-Adapter (Ye et al., 2023) with its decoupled cross-attention design and technique like PhotoMaker (Li et al., 2024) that accelerates generation using identity images. Recently, StoryDiffusion (Zhou et al., 2024) introduced Consistent Self-Attention (CSA) to boost the frame-wise subject consistency but still faces limitations in long image sequences. In contrast, Story-Adapter maintains image semantic consistency in long image sequences by using cross-attention on global embeddings from all generated images of the previous iteration and the corresponding text features. Along with our iteration paradigm, the whole generations are gradually improved w.r.t semantic consistency and generative quality for fine-grained interactions.
@@ 3 Method
Figure 3: Illustration of the proposed iterative paradigm, which consists of initialization, iterations in Story-Adapter, and implementation of Global Reference Cross-Attention.
Compared to regular-length stories, long stories contain more characters and more complex interactions, leading to higher requirements for semantic consistency and fine-grained interaction generation. To address the above challenges, we resort to an iterative paradigm that progressively refines all the generated images, w.r.t. semantic consistency and visual details in multiple rounds. We instantiate the iterative paradigm by equipping a fixed Stable-Diffusion (SD) model with a cross-attention mechanism, termed Story-Adapter. The pipeline is demonstrated in Fig. 3.
@@ 3.1 Initialization
To build the initialization for iteration, we only employ text prompt $T_{k}$ for the $k_{th}$ image in the story to guide the fixed ${\rm{SD}}(z,T_{k})$ in generating the initial images, where $z$ is the random noise. All generated images from the initial step will be stored as reference images for the first iteration. We denote $i=0$ as the initialization of Story-Adapter. Thus, the whole initialization process can be represented as:
$\begin{split}&{x_{k}^{i=0}}={\rm{SD}}(z,T_{k}),k\in[1,B],\\ &{x_{1,\cdots,B}^{i=0}}=[x_{1}^{i=0},x_{2}^{i=0},\cdots,x_{k}^{i=0},\cdots,x_{B-1}^{i=0},x_{B}^{i=0}],\end{split}$ (1)
@@
where $B$ denotes the length of the story. Compared to subject-consistent image generation methods (Ye et al., 2023) that introduce reference image guidance, initialization which relies only on text prompts more faithfully visualizes the corresponding content in the story. The following iterations benefit from the rich visual content provided by the initialization of the reference images.
3.2 Story-Adapter
@@
This subsection demonstrates how each image is updated within an iteration in Story-Adapter. Formally, for the $i_{th}$ iteration, we use all visualizations from the previous iteration $x_{1,\cdots,B}^{i-1}$ as the reference images $R$ to refine the generated images in the current round. For the generation of the $k_{th}$ image of a long story, we define a function ${\rm{SD_{GRCA}}}(z,T_{k},x_{1,\cdots,B}^{i-1})$ to represent the whole denoising process with our Global Reference Cross-Attention (Sec. 3.3).
Thus the $i_{th}$ iteration can be expressed as:
$\begin{split}&{x_{k}^{i}}={\rm{SD_{GRCA}}}(z,T_{k},x_{1,\cdots,B}^{i-1}),k\in[1,B],\\ &{x_{1,\cdots,B}^{i}}=[x_{1}^{i},x_{2}^{i},\cdots,x_{k}^{i},\cdots,x_{B-1}^{i},x_{B}^{i}],\\ \end{split}$ (2)
As iterations proceed, the reference images evolve to be more coherent, as Story-Adapter consistently improves the semantic consistency in a global view. Additionally, generative quality for fine-grained interactions is also constantly optimized as Story-Adapter repeatedly engages text prompt constraints during iterations.
Algorithm 1 Pseudo-Code of Story-Adapter.
1 # diffusion model:$\theta$, iteration epochs:$\rm{L}$, starting weight factor:$\lambda_{s}$, ending weight factor:$\lambda_{e}$, $i_{th}$ iteration $j_{th}$ diffusion step $k_{th}$ intermediate denoising features:$I_{k,j}^{i}$, story length:$\rm{B}$, diffusion steps:$\rm{J}$, decoder:$\rm{D}$
2 # Initialize $I_{k,j}^{0}$, $I_{k,j}^{i}$~N(0, I), k~(1, B), i~(1, L), j~(0, J)
3 # Initialize Story-Adapter iteration
4 for j in reversed(range(0, J)):
5 # Init z~N(0, I) if j>1 else z=0
6 $I_{k,j-1}^{0}$=(1/sqrt($\alpha_{j}$))*$I_{k,j}^{0}$-(1-$\alpha_{j}$)*$\theta$($I_{k,j}^{0}$,j,$T_{k}$)/sqrt(1-$\alpha_{j}$))+$\sigma_{t}$*z
7 R=concat([$x_{1}^{0}$,...,$x_{k}^{0}$,...,$x_{B}^{0}$]), $x_{k}^{0}$=D($I_{k,0}^{0}$)
8
9 # Insert GRCA to $\theta$ and initialize weighting factor list $\lambda_{list}$
10 $\lambda_{list}$=linspace($\lambda_{s}$,$\lambda_{e}$,L)
11 # Story-Adapter Iteration
12 for i,$\lambda$ in enumerate($\lambda_{list}$):
13 for j in reversed(range(0, J)):
14 $I_{k,j-1}^{i}$=(1/sqrt($\alpha_{j}$))*($I_{k,j}^{i}$-(1-$\alpha_{j}$)*$\theta$($I_{k,j}^{i}$,j,$T_{k}$,R,$\lambda$)/sqrt(1-$\alpha_{j}$))+$\sigma_{t}$*z
15 R=concat([$x_{1}^{i}$,...,$x_{k}^{i}$,...,$x_{B}^{i}$]), $x_{k}^{i}$=D($I_{k,0}^{i}$)
@@ 3.3 Global Reference Cross-Attention
Although incorporating image context or reference images extends text-to-image generation to character-consistent image sequences, existing AR paradigms (Pan et al., 2024; Liu et al., 2024) suffer from error accumulation over long stories, while RI paradigms (Zhou et al., 2024) may propagate flaws from the reference images.
@@
In contrast, we propose an efficient plug-and-play augmentation module to equip SD models, called Global Reference Cross-Attention (GRCA). We utilize a pre-trained CLIP (Radford et al., 2021) image encoder to extract a global embedding $c$ for each reference image from the previous round, effectively preserving the semantics of reference images using only a few tokens. The token simplification allows GRCA to incorporate all reference images as guidance in the cross-attention process without incurring significant computational overhead.
@@
In the $i_{th}$ iteration of Story-Adapter, given all the reference images in the previous round $x_{1,\cdots,B}^{i-1}\in{\mathbb{R}^{{B}\times{h}\times{w}\times\rm{3}}}$, ${h}$, ${w}$ denote reference image resolution. We define a function ${\rm{Attention}}(Q,K,V)$ to indicate the attention calculation, where $Q$, $K$, and $V$ represent the query, key, and value in the attention, respectively. GRCA in the visualization for the $k_{th}$ image can be specified as:
@@
$\begin{split}&{c_{1,\cdots,B}^{i}}={\rm{CLIP}}(x_{1,\cdots,B}^{i-1}),{c_{1,\cdots,B}^{i}}\in{\mathbb{R}^{{B}\times{d}}},\\ &{c_{1,\cdots,B}^{i}}={c_{1,\cdots,B}^{i}}{W_{c}},{W_{c}}\in{\mathbb{R}^{{d}\times{ne}}},\\ &{c_{1,\cdots,B}^{i}}={\rm{flatten}}({c_{1,\cdots,B}^{i}}),{c_{1,\cdots,B}^{i}}\in{\mathbb{R}^{{1}\times{Bn}\times{e}}},\\ &{{Q}_{k}^{i}}={{I_{k}}{W_{q}}},{{K_{k}^{i}}}={{c_{1,\cdots,B}^{i}}{W_{k}}},{{V}_{k}^{i}}={{c_{1,\cdots,B}^{i}}{W_{v}}},\\ &{{\rm{GRCA}}({I_{k}},{x_{1,\cdots,B}^{i-1}})}={{\rm{Attention}}({{Q}_{k}^{i}},{{K_{k}^{i}}},{{V}_{k}^{i}})},\end{split}$ (3)
@@
Where $W_{c}$ is the projection matrix of global embeddings transformed into reference tokens. $d$ and $e$ denote the embedded dimension of global embeddings and the projection dimension of the projection matrix, respectively. $n$ indicates the number of reference tokens for a single reference image, $n=4$ if not specified. $\rm{flatten(.)}$ represents a flatten operation for vectors. $W_{q}$ is the mapping weight matrix for the intermediate denoising feature $I$ in SD. $W_{k}$, $W_{v}$ are the mapping weight matrices of the reference tokens.
@@
Eventually, we merge the outputs from GRCA with the outputs from text cross-attention, to guide the visualization of $k_{th}$ image in the story. In particular, with corresponding text prompt $T_{k}$ and all reference images $x_{1,\cdots,B}^{i-1}$, the intermediate denoising feature ${I}_{k}^{i}$ is obtained as follows:
$\begin{split}&{{I}_{k}^{i}}={{\rm{Attention}}({I}_{k}^{i},{T}_{k},{T}_{k})}+\lambda{{\rm{GRCA}}({I_{k}^{i}},{x_{1,\cdots,B}^{i-1}})}.\\ \end{split}$ (4)
where $\rm{\lambda}$ is a balance factor for controlling the influence of GRCA on the visualization results. We propose a linear weighting strategy to adjust the weight factor for each iteration, where the weight factor increases linearly with a low value to trade off visual consistency and text alignment in the iterative paradigm. Since the existing diffusion models contain a cross-attention design associated with the reference image, our GRCA could be directly plugged in and reuse the cross-attention weights without training. We demonstrate the procedure of Story-Adapter, along with the linear weighting strategy in Algo. 1.

### v2

# Story-Iter: A Training-free Iterative Paradigm for Long Story Visualization

@@ Abstract
This paper introduces Story-Iter, a new training-free iterative paradigm to enhance long-story generation. Unlike existing methods that rely on fixed reference images to construct a complete story, our approach features a novel external iterative paradigm, extending beyond the internal iterative denoising steps of diffusion models, to continuously refine each generated image by incorporating all reference images from the previous round. To achieve this, we propose a plug-and-play, training-free global reference cross-attention (GRCA) module, modeling all reference frames with global embeddings, ensuring semantic consistency in long sequences. By progressively incorporating holistic visual context and text constraints, our iterative paradigm enables precise generation with fine-grained interactions, optimizing the story visualization step-by-step. Extensive experiments in the official story visualization dataset and our long story benchmark demonstrate that Story-Iter’s state-of-the-art performance in long-story visualization (up to 100 frames) excels in both semantic consistency and fine-grained interactions.
Figure 1: A long story of “snowman” visualized by our Story-Iter from different iterations, compared with those visualized by previous StoryDiffusion Zhou et al. (2024) and StoryGen Liu et al. (2024). Notable differences are highlighted in green and red. Zoom in for a better view.
Figure 2: Comparison of paradigms for long story visualization: (A) Auto-Regressive (AR): generates frames sequentially referencing previous finite frames (e.g. the previous three frames); (B) Reference-Image (RI): employs fixed reference images (e.g. the beginning four frames) as reference images; (C) Iterative Paradigm: leverages all frames from the previous iteration as reference images.
@@ 1 Introduction
Story visualization aims to generate a sequence of coherent images from text prompts, reflecting the narrative’s progression and enabling users, even without an artistic background, to visually present their stories (Li et al., 2019; Maharana and Bansal, 2021; Chen et al., 2022). Recent advancements in text-to-image models, particularly diffusion models, have significantly improved the quality of generated visuals, producing high-quality, creative, and aesthetically pleasing images (Saharia et al., 2022; Rombach et al., 2022; Kang et al., 2023). These models greatly outperform earlier approaches, such as generative adversarial networks (Brock, 2018) in terms of image quality.
@@
However, story visualization remains challenging, particularly in maintaining semantic consistency and synthesizing interactions as the story length increases. Two main paradigms have emerged in this domain. The Auto-Regressive Paradigm (Fig. 2 (A)), which generates frames sequentially referencing previous finite frames (Pan et al., 2024; Liu et al., 2024), often struggles with semantic consistency due to error accumulation and the inability to reference future frames. Although techniques like Consistent Self-Attention (CSA) (Zhou et al., 2024) can help mitigate these inconsistencies, their reliance on intermediate denoising features results in high memory consumption, limiting scalability for longer stories. To address these challenges, Zhou et al. (2024) further proposes the Reference-Image Paradigm (Fig. 2 (B)), which employs fixed reference images to guide the visualization. However, while using only the initial frames as reference images alleviates scalability issues, it fails to provide the global semantic coherence necessary for long-story visualization, resulting in error propagation from the reference images to subsequent frames. As such, both paradigms experience quality degradation when visualizing long stories. Additionally, they inherit the limitations from Stable Diffusion Model (SDM) (Rombach et al., 2022), particularly in generating fine-grained interactions in the story (Fig. 1).
To address the limitations in existing story generation methods, we present Story-Iter, a new Iterative Paradigm. Unlike existing methods that use multiple fixed reference frames but do not refine them over time or incorporate a holistic visual context (Fig. 2 A&B, e.g. StoryGen (Liu et al., 2024) and StoryDiffusion (Zhou et al., 2024)), our approach continuously refines each generated story frame, utilizing the full-length reference images generated from the previous iteration. Here, iteration refers to the external round beyond the internal denoising steps of diffusion models.
The proposed iterative paradigm offers two key advantages. 1) It progressively approximates the distribution of reference image global embeddings across iterations, ultimately ensuring visual coherence (see Fig. 4). 2) By iteratively engaging the full-length story frames and text prompts, it optimizes fine-grained control based on both global and local semantic contexts. Shown in Fig. 1, Story-Iter enhances both visual consistency and fine control across iterations, resulting in more coherent and higher-quality visualizations. For example, the image depicting complex character interactions, such as “snowman saw a fox” demonstrates substantial improvement over iterations compared to previous methods (Liu et al., 2024; Zhou et al., 2024).
Figure 3: Attention maps for each generated image in our Global Reference Cross-Attention. Zoom in for a better view.
Figure 4: Distribution among the global embeddings of full-length reference images as iterations progress.
@@
Specifically, during initialization, only text prompts of the story are utilized to generate the full story frames without any reference. In the subsequent iterations, the global embeddings of the full-length story frames from the previous iteration as the reference, along with the text embeddings, collaboratively guide story generation. To implement the iterative paradigm, we introduce a plug-and-play Global Reference Cross-Attention (GRCA) for long story visualization. Some subject-consistent generation methods, such as the decoupled cross-attention in IP-Adapter (Ye et al., 2023), also contribute to story visualization. However, they are designed to maintain semantic consistency in individual images, whereas story visualization requires connections across all frames, making them not directly applicable. Different from the IP-Adapter, the global embeddings of the full-length reference images act as keys and values in GRCA. Unlike the intermediate denoising features used in CSA in StoryDiffusion (Zhou et al., 2024), the global embeddings in GRCA allow more image frames to be referenced to maintain global semantic consistency. The processes of computing the token similarity matrix and merging tokens in GRCA adaptively aggregate visual features from reference images similar to the currently generated image (Fig. 4). This ultimately ensures the semantic consistency in the iterative paradigm and reduces the influence from noisy references. Additionally, to strike a balance between visual consistency and text controllability, we introduce a linear weighting strategy to fuse both modalities.
@@
Extensive experiments demonstrate that Story-Iter consistently outperforms existing methods for visualizing both regular-length and long stories (up to 100 frames). Specifically, in the context of regular-length story visualization using the StorySalon benchmark (Liu et al., 2024), Story-Iter exceeds the baseline model, StoryGen (Liu et al., 2024), achieving a 9.4% improvement in average Character-Character Similarity (aCCS) (Cheng et al., 2024b) and a 21.71 reduction in average Fréchet Inception Distance (aFID) (Cheng et al., 2024b). For long story visualization, Story-Iter also demonstrates solid advancements, achieving gains of 3.4% in aCCS and 8.14 in aFID compared to StoryDiffusion (Zhou et al., 2024), demonstrating the superior generative quality of Story-Iter, particularly in semantic consistency and fine-grained interactions.
Our contributions are summarized as follows:
  • A new long story benchmark for evaluating long story visualization for up to 100 frames.
  • A new iterative paradigm that enhances story consistency by continuously updating reference images in each external iteration, beyond the diffusion model’s internal denoising steps.
  • A new global attention mechanism GRCA that enables modeling all frames as reference images, thus ensuring semantic consistency in long sequences.
  • State-of-the-art story visualization performance on various benchmarks, especially in long-story scenarios.
@@ 2 Related Work
Story Visualization
Story visualization (Chen et al., 2022; Li, 2022; Tao et al., 2025; Papadimitriou et al., 2024; Bugliarello et al., 2024; Ahn et al., 2023; Rahman et al., 2023; Tsakas et al., 2023; Wu et al., 2024) has evolved from GAN-based approaches like StoryGAN (Li et al., 2019) to more advanced techniques. Recent developments leverage diffusion models (Shen et al., 2024; Tao et al., 2024) and combine them with the auto-regressive paradigm, as seen in AR-LDM (Pan et al., 2024) and StoryGen (Liu et al., 2024). StoryDiffusion (Zhou et al., 2024), on the other hand, uses a reference-image paradigm with fixed reference images. DreamStory (He et al., 2024), StoryGPT (Shen and Elhoseiny, 2023), and MovieDreamer (Zhao et al., 2024) introduce LLM-generated guidance for story visualization. There are also attempts on interactive story visualization (Cheng et al., 2024a; b; Gong et al., 2023; Yang et al., 2024). However, challenges remain in maintaining semantic consistency for the whole story and avoiding error accumulation, especially for longer narratives (Wang et al., 2023; Tewel et al., 2024; Zhou et al., 2024; Liu et al., 2024). Unlike StoryDiffusion (Zhou et al., 2024), which uses fixed reference frames but does not refine them or incorporate holistic visual context, our paradigm continuously refines each generated frame, utilizing the full-length story frames generated from the previous external iteration.
@@ 3.1 Initialization
To build the initialization for iteration, we only employ text prompt $T_{k}$ for the $k_{th}$ image in the story to guide the pretrained Stable Diffusion Model (Rombach et al., 2022) ${\rm{SDM}}(z,\,T_{k})$ in generating the initial images, where $z$ is the random noise. All generated images from the initial step will be stored as reference images for the first iteration. We denote $i=0$ as the initialization of Story-Iter. Thus, the initially generated story frames can be represented as:
$\begin{split}&{x_{k}^{i=0}}={\rm{SDM}}(z,\,T_{k}),\,k\in[1,\,B],\\ &{x_{1\cdots B}^{i=0}}=[x_{1}^{i=0},\,x_{2}^{i=0},\cdots,x_{k}^{i=0},\cdots,\,x_{B-1}^{i=0},\,x_{B}^{i=0}],\end{split}$ (1)
@@
where $B$ denotes the story length. Compared to subject-consistent image generation methods (Ye et al., 2023) that introduce reference image guidance, Story-Iter is initialized only by text prompts to outline the story faithfully.
Figure 5: Overview of Story-Iter. The full-length story frames are generated with only text prompts during initialization (Sec. 3.1). In the subsequent external iterations, the story frames are generated referencing the full-length story frames from the previous iteration (Sec. 3.2). The Global Reference Cross-Attention (GRCA) adaptively aggregates features from full-length reference images to ensure semantic consistency (Sec. 3.3).
3.2 Iterative Paradigm
@@
In this subsection, we describe how each image is updated within an external iteration beyond the internal denoising steps of diffusion models (see Fig. 5). Formally, for the $i_{th}$ iteration, we use the full-length $B$ story frames from the $(i-1)_{th}$ iteration $x_{1\cdots B}^{i-1}$ ($i\in[1,\,L]$, $L$ refers to total iteration number) as the reference images to refine story visualization. With ${\rm{SDM_{GRCA}}}(z,\,T_{k},\,x_{1\cdots B}^{i-1})$ representing the whole denoising process with our Global Reference Cross-Attention in pretrained Stable Diffusion Model (Sec. 3.3), the $k_{th}$ story frame generated from the $i_{th}$ iteration can be expressed as:
[…]
As the iterations progress, the distribution of reference image global embeddings gradually converges, thereby assisting Story-Iter in continuously enhancing the semantic consistency of the global view. As shown in Fig. 4, Story-Iter does not enforce consistency in a single step. Instead, through successive external iterations, it progressively refines GRCA by aligning the current frame’s features with semantically relevant content from reference frames. Fine-grained interactions (limitation for SDM) are also optimized as Story-Iter repeatedly engages text constraints $T_{k}$ in the iterative paradigm.
@@ 3.3 Global Reference Cross-Attention
We propose a plug-and-play augmentation module to equip Stable Diffusion models, called Global Reference Cross-Attention (GRCA). We utilize pre-trained CLIP (Radford et al., 2021) as the image encoder and project the global embedding (termed $c$) for each reference image into a few tokens. This global embedding enables guidance of structural and visual similarity to reference images during the denoising process. Furthermore, with fewer tokens, GRCA can incorporate more reference images.
@@
Given full-length $B$ images generated from the $(i-1)_{th}$ iteration $x_{1\cdots B}^{i-1}\in{\mathbb{R}^{{B}\times{h}\times{w}\times\rm{3}}}$ (h{h}: height, w{w}: width), we define ${\rm{Attention}}(Q,K,V)$ to indicate the attention calculation, where $Q$, $K$, and $V$ represent the query, key, and value respectively. GRCA of the $k_{th}$ image in the $i_{th}$ iteration:
@@
$\begin{split}&{c_{1\cdots B}^{i}}={\rm{CLIP}}(x_{1\cdots B}^{i-1}),\quad{c_{1\cdots B}^{i}}\in{\mathbb{R}^{{B}\times{d}}},\\ &{\hat{c}_{1\cdots B}^{i}}={c_{1\cdots B}^{i}}{W_{c}},\quad{W_{c}}\in{\mathbb{R}^{{d}\times{(n\times e)}}},\\ &{\tilde{c}_{1\cdots B}^{i}}={\rm{flatten}}({\hat{c}_{1\cdots B}^{i}}),\quad{\tilde{c}_{1\cdots B}^{i}}\in{\mathbb{R}^{{1}\times{(B\times n)}\times{e}}},\\ &{{Q}_{k}^{i}}={{I_{k}}{W_{q}}},\,{{K_{k}^{i}}}={{\tilde{c}_{1\cdots B}^{i}}{W_{k}}},\,{{V}_{k}^{i}}={{\tilde{c}_{1\cdots B}^{i}}{W_{v}}},\\ &{{\rm{GRCA}}({I_{k}^{i}},\,{x_{1\cdots B}^{i-1}})}={{\rm{Attention}}({{Q}_{k}^{i}},\,{{K_{k}^{i}}},\,{{V}_{k}^{i}})}.\end{split}$ (3)
@@
The dimension of global embedding $c_{k}^{i}$ and the projected dimension are denoted by $d$ and $e$, respectively. $n$ indicates the length of reference tokens, $n=4$ if not specified. $\rm{flatten(.)}$ represents a flatten operation on vectors. $W_{c}$ is the projection matrix transforming the global embeddings into reference tokens. $W_{q}$ is the weight matrix of the intermediate denoising feature $I_{k}^{i}$ in SDM. $W_{k}$ and $W_{v}$ are the weight matrices of the reference tokens. The proposed plug-and-play GRCA directly reuses the cross-attention weights from IP-Adapter (Ye et al., 2023) in subject-consistent single-image generation without training.
Figure 7: Qualitative comparisons for long story visualization. Story-Iter shows advantages in generating semantic consistency (consistency and differences are highlighted in green and red boxes) and character interactions (correct and incorrect interactions are highlighted in green and red circles). For the complete story, refer to the Fig. 24.
@@
Finally, we merge the outputs from GRCA with the outputs from cross-attention on texts, to guide the $k_{th}$ story frame generation. With corresponding text prompt $T_{k}$ and the full-length reference images $x_{1\cdots B}^{i-1}$, the intermediate denoising feature ${I}_{k}^{i}$ is updated as:
[…]
where the weight factor $\lambda_{i}$ of the $i_{th}$ iteration is adjusted as $\lambda_{i}=\lambda_{1}+q\times(i-1)$ with $q$ fixed. $\lambda_{i}$ increases linearly with a lower bound $\lambda_{1}$ to balance visual consistency and text alignment in the iterative paradigm. Leveraging pretrained weights, GRCA adaptively re-weights attention at each iteration, thereby mitigating the accumulation of noise from irrelevant historical states and instead prioritizing information most pertinent to the current frame (see Fig. 4). The global embedding undergoes progressive updates, serving as a soft forgetting mechanism that attenuates contextual drift and naturally suppresses obsolete or extraneous information throughout the iterative refinement process. We demonstrate Pseudo-code of Story-Iter in Algo. 1.